### v1

# Availability-aware Sensor Fusion via Unified Canonical Space for 4D Radar, LiDAR, and Camera

Abstract
Sensor fusion of camera, LiDAR, and 4-dimensional (4D) Radar has brought a significant performance improvement in autonomous driving (AD). However, there still exist fundamental challenges: deeply coupled fusion methods assume continuous sensor availability, making them vulnerable to sensor degradation and failure, whereas sensor-wise cross-attention fusion methods struggle with computational cost and unified feature representation. This paper presents availability-aware sensor fusion (ASF), a novel method that employs unified canonical projection (UCP) to enable consistency in all sensor features for fusion and cross-attention across sensors along patches (CASAP) to enhance robustness of sensor fusion against sensor degradation and failure. As a result, the proposed ASF shows a superior object detection performance to the existing state-of-the-art fusion methods under various weather and sensor degradation (or failure) conditions; Extensive experiments on the K-Radar dataset demonstrate that ASF achieves improvements of 9.7% in $AP_{BEV}$ (87.2%) and 20.1% in $AP_{3D}$ (73.6%) in object detection at IoU=0.5, while requiring a low computational cost. The code will be available at https://github.com/kaist-avelab/K-Radar.
Figure 1: Visualization of feature representation with t-SNE [1] at different stages of ASF for ‘Sedan’ class. Red, green, blue, and gray dots represent features from camera, LiDAR, 4D Radar, and fused features, respectively. Symbols in solid lines such as circle and triangle, square, and star indicate normal, sleet, and heavy snow conditions, respectively. (a) Initial output features from sensor-specific encoders show inconsistent distribution across sensors. (b) After UCP, features become better aligned to the fused feature. (c) After CASAP-PN, features from available sensors form cohesive clusters (in dashed symbols) based on weather conditions. Note that in adverse weather condition, camera features show larger deviation due to the degradation. Additional t-SNE visualizations are in Appendix B.
Figure 2: Comparison of sensor fusion methods: (a) DCF (e.g., 3D-LRF [2]), (b) SCF (e.g., DPFT [3]), and (c) ASF. FV, Obj., GT, $\mathcal{L}_{cls}$ and $\mathcal{L}_{reg}$ stand for ‘front-view’, ‘objects’, ‘ground truths’, ‘classification loss’, and ‘regression loss’, respectively. ‘Feature coupling’ refers to methods that combine features from multiple sensors to create new features. Optional components are in dashed lines; for example, [4] combines camera and LiDAR features without transforming the camera viewpoint, while [5] fuses features through a transformer decoder head [6] without explicit feature coupling. ASF does not apply feature coupling to ensure independence between sensors.
1 Introduction
Autonomous driving (AD) technology has advanced rapidly, with multiple companies adopting multi-sensor fusion approaches that combine two or more sensors, such as cameras, LiDAR, and 4-dimensional (4D) Radar, to achieve more robust and reliable perception [7, 8, 9]. Cameras uniquely provide color information but struggle with depth estimation; LiDAR delivers high-resolution 3-dimensional (3D) point cloud data but is less reliable in adverse weather conditions [10]; and 4D Radar, despite its relatively low angular resolution, offers robustness in adverse weather and directly measures relative velocity [11, 12, 13, 14, 15, 16]. This complementary nature initiated sensor fusion between camera, LiDAR, and 4D Radar to improve perception performance and reliability compared to a single-sensor [5, 17, 18, 19, 20].
Most multi-modal sensor fusion methods can be categorized into two methods. The first is deeply coupled fusion (DCF), which directly combines feature maps (FMs) extracted by sensor-specific encoder, as illustrated in Fig. 2-(a). While it is simple to implement and computationally efficient with excellent performance in various benchmarks [2, 18, 17, 21, 22, 23], it assumes all sensors are functioning properly and consistently. This makes DCF vulnerable to sensor degradation due to adverse weathers, surface-damages, and sensor failure. Moreover, DCF requires retraining the entire neural network when the number of sensors changes, as the size of the fused FM (i.e., the input to the detection head) changes. The second method is sensor-wise cross-attention fusion (SCF), which divides features extracted from each sensor into patches with positional-embedding (e.g., depth information for camera [24, 25]) and selectively combines available patches using cross-attention (Fig. 2-(b)), allowing it to handle cases where some sensors are degraded. However, SCF does not have sensor scalability [26], since the method does not project sensor-specific features into a standardized representation [27, 28, 29, 30, 31], and SCF incurs computational complexity that increases quadratically with the number of patches [3, 5, 32].
One of the fundamental limitations of existing fusion methods stems from inconsistencies in feature representation across different sensors [26, 33]. Cameras produce 2D RGB images, whereas LiDAR generates 3D point clouds, and 4D Radar produces low-resolution tensors with power values. Therefore, each sensor extracts features with different representations for the same object, making consistent fusion challenging (as shown in Fig. 1-(a)). To address this inconsistency, an ideal strategy could project features from different sensors into a unified canonical space for fusion. The concept of ‘True Redundancy [26]’, that ensures sensor independence while maintaining canonical feature representation for any sensor combination, suggests a promising direction for highly reliable and robust sensor fusion.
Therefore, we propose availability-aware sensor fusion (ASF) method (Fig. 2-(c)), in which each sensor (i.e., camera, Lidar, and 4D Radar) performs independently while being complementarily fused through a projection to a unified canonical space. As a result, the proposed method addresses the limitations of both DCF and SCF simultaneously. The key innovation of ASF is in two sub-modules; First, unified canonical projection (UCP) projects features from each sensor into a unified space based on common criteria (i.e., canonical). Since UCP is optimized using the same reference query for all sensors, inconsistencies in sensor features are eliminated. While Fig. 1-(a) shows sensor features represented without clear patterns, Fig. 1-(b) demonstrates how UCP aligns the features from each sensor to the fused feature. Second, cross-attention across sensors along patches with post-feature normalization (CASAP-PN) estimates the availability of sensors through patch-wise cross-attention on features projected into the unified canonical space, assigning higher weights to features from available sensors and lower weights to features from missing or uncertain sensors. Unlike SCF that applies the cross-attention across all sensors ($N_{s}$) and patches ($N_{p}$) (i.e., $O(N_{s}^{2}N_{p}^{2})$), ASF only applies the cross-attention across sensors along patches (i.e., $O(N_{s}^{2}N_{p})$). Because of this, ASF eliminates complex positional-embedding and improves computationally efficiency. Additionally, it applies normalization to ensure that (camera, LiDAR, and 4D Radar) features can be processed consistently by the detection head regardless of sensor combination. In contrast to Fig. 1-(b), where each sensor’s features remain separate, Fig. 1-(c) shows how sensor features cluster together after post-feature normalization (except for camera images become useless in adverse weather conditions, as shown in Fig. 3). This enables ASF to flexibly handle sensor degradation or failure and to achieve the reliability embodied in the ‘True Redundancy’ concept [26] for AD implementation.
To integrate the availability awareness into the detection head, we propose a sensor combination loss (SCL) that optimizes learning outcomes across all sensor combinations. SCL considers individual sensor unavailability during the training, enabling the system to maintain high performance in the presence of unexpected sensor failure or adverse weather conditions. The effectiveness of our proposed ASF method has been validated on the K-Radar dataset [12], demonstrating improvements of 9.7% in $AP_{BEV}$ (87.2%) and 20.1% in $AP_{3D}$ (73.6%) for detection performance at IoU=0.5 compared to state-of-the-art (SOTA) fusion methods [2, 34], which includes the performance in extreme situations such as sensor degradation or failure (unavailable).
The main contributions of this paper are as follows:
  • We propose ASF based on UCP and CASAP-PN that achieves superior performance to SOTA methods and robust performance against sensor degradation and failure.
  • We propose SCL for the loss function to optimize the detection performance for all possible sensor combinations.
  • Through extensive experiments on the K-Radar dataset, we demonstrate that ASF achieves the high performance with low computational load.
The remainder of this paper is organized as follows: Section 2 introduces existing methods, and Section 3 describes the proposed ASF in detail. Section 4 presents experimental settings and results on the K-Radar dataset, analyzing performance in various scenarios including sensor degradation and failure. Section 5 concludes the paper with a summary and discusses the limitations. All code and logs for ASF will be made available at the K-Radar github repository.
2 Related Works
2.1 Deeply Coupled Fusion (DCF)
DCF constructs a fused feature map (FM) by concatenating FMs from each sensor. Most studies focus on fusing camera with LiDAR or 4D Radar [18, 17, 19, 20], or combining LiDAR and 4D Radar [2, 34]. Implementations range from directly fusing the front-view (FV) camera image with LiDAR or 4D Radar without view transformation [4] to applying learnable BEV transforms [35] and concatenating at the BEV stage [18, 17, 19]. DCF method is straightforward to implement and computationally efficient compared to SCF, demonstrating strong performance across multiple benchmarks.
Recent DCF studies have improved performance by applying feature coupling, where the FMs of each sensor are enhanced with FMs of other sensors using multilayer perception (MLP) or attention mechanisms [36]. 3D-LRF [2] demonstrated superior performance to the conventional DCF methods [37, 17, 18] by applying attention between LiDAR and 4D Radar voxel features before concatenation. L4DR [34] achieved SOTA performance on the K-Radar dataset by weighting each sensor’s FM using coupled BEV FMs with LiDAR and 4D Radar. However, DCF assumes that all sensors are functioning normally, as they rely on the fused FM constructed by concatenating FMs from all sensors, which makes DCF performance vulnerable to sensor degradation and failure.
2.2 Sensor-wise Cross-attention Fusion (SCF)
SCF divides FMs of all sensors into patches and dynamically combines patches of all sensors through cross-attention in a transformer decoder head [3, 32, 5]. This method can inherently utilize sensor availability through the cross-attention. TransFusion [32] is the first SCF method that addresses sensor availability, but its sequential fusion method (e.g., performing LiDAR detection first and then fusing with camera data) makes inference impossible when LiDAR is unavailable. CMT [5] presents an availability-aware sensor fusion of camera and LiDAR data using a transformer decoder head without applying feature coupling to individual sensors. However, without feature coupling for camera, CMT relies on positional embedding to incorporate depth information into camera FMs [24], which results in 3D patches ($H\!\times\!W\!\times\!D$). This leads to computational complexity of $O(N_{s}^{2}N_{p}^{2})$ (where $N_{s}$ and $N_{p}$ are the number of sensors and patches, respectively), causing explosive growth in computational cost and memory usage. For instance, CMT requires 8 A100 GPUs with 80GB VRAM to train with a batch size of 16. Recently, DPFT [3] projects depth-aware 4D Radar FMs onto the FM from the front-view camera image to create patches and verify sensor availability. However, it achieves relatively low performance (56.1% $AP_{3D}$ at IoU=0.3) but demands high computational cost that scales quadratically with the number of patches.
In summary, SCF methods require high computational cost and directly utilize FM from each sensor without a projection process, making it difficult to create canonical FMs in a unified space and to improve sensor scalability. This makes SCF methods challenging to implement in resource-constrained real-time AD systems.
3 Proposed Methods
3.1 Sensor Fusion Framework
The overall sensor fusion framework consists of three stages: (1) sensor-specific encoders (i.e., backbones) that extract same-sized bird’s eye-view (BEV) feature maps (FMs) from each sensor data (i.e., RGB image, LiDAR point cloud, and 4D Radar tensor), (2) the proposed availability-aware sensor fusion (ASF) network that is described in following subsection 3.2, and (3) a detection head that detects objects from the fused FM.
Focusing on our ASF contribution, we utilize established methods for the sensor-specific encoders and detection head. Specifically, we adopt BEVDepth [38], SECOND [39], and RTNH [12] backbones for camera, LiDAR, and 4D Radar, respectively, along with a SSD detection head [40, 41]. Further specifications regarding the overall sensor fusion framework, such as sensor-specific encoders and detection head, are provided in Appendix A.
3.2 Availability-aware Sensor Fusion (ASF)
As illustrated in Fig. 2-(c), the proposed ASF consists of two key components: unified canonical projection (UCP) and cross-attention across sensors along patches with post-feature normalization (CASAP-PN).
Unified Canonical Projection (UCP). One of the key challenge in multi-modal sensor fusion is the inherent inconsistency of features from different sensors [26, 33], as visualized in Fig. 1-(a). To tackle this, we divide BEV FMs into an equal number of patches for all sensors and train projection functions to transform features from each sensor into a unified space based on the same criteria (i.e., reference query in CASAP-PN). To formally define our methods, we first represent the same-sized FMs of each sensor as:
$\mathrm{\mathbf{FM}}^{s}\in\mathbb{R}^{C_{s}{\times}H{\times}W},s\in\{S_{C},S_{L},S_{R}\},$ (1)
where $C_{s}$ denotes the channel dimension for sensor $s$, $H$ and $W$ represent the identical height and width of the BEV FMs for all sensors, respectively, and $S_{C}$, $S_{L}$, and $S_{R}$ denote camera, LiDAR, and 4D Radar sensors, respectively. Each FM is then divided into patches $\mathbf{F}_{p}^{s}$ with height $P_{H}$ and width $P_{W}$:
$\mathcal{T}_{p}(\mathbf{FM}^{s})\!=\!\{\mathbf{F}_{p,i}^{s}|\mathbf{F}_{p,i}^{s}\!\in\!\mathbb{R}^{C_{s}\times P_{H}\times P_{W}},i\!=\!1\!:\!N_{p}\},$ (2)
where $\mathcal{T}_{p}(\cdot)$ is the operation that divides each FM into patches, $N_{p}\!=\!(H/P_{H})\!\times\!(W/P_{W})$ is the number of patches, which is identical across all sensors since each FM has the same spatial size, and ‘$1\!:\!N_{p}$’ denotes ‘$1,2,...,N_{p}$’. Note that since the patches are already spatially aligned (i.e., $\mathbf{F}_{p,i}^{S_{C}}$, $\mathbf{F}_{p,i}^{S_{L}}$, and $\mathbf{F}_{p,i}^{S_{R}}$ correspond to the same position), our method eliminates the use of computationally expensive positional-embedding [24] required for SCF [5, 3]. Then, we apply a parallel operation along patches that projects each patch to have the same channel dimension $C_{u}$. This is the UCP operation $\mathcal{U}^{s}(\cdot)$ that transforms sensor-specific patches into patches in a unified canonical space. The UCP-processed patch $\mathcal{P}_{u}^{s}$ for each sensor is expressed as follows:
$\mathcal{P}^{s}_{u}\!=\!\{\mathbf{F}_{u,i}^{s}|\mathbf{F}_{u,i}^{s}\!=\!\mathcal{U}^{s}(\mathrm{LN}(\mathbf{F}_{p,i}^{s}))\!\in\!\mathbb{R}^{C_{u}},i\!=\!1\!:\!N_{p}\}$ (3)
$\mathcal{U}^{s}(\cdot)\!=\!\mathrm{LN}(\mathrm{Proj^{(n_{u})}(\cdot)}),\mathrm{Proj(\cdot)\!=\!\mathrm{GeLU}(\mathrm{MLP}(\cdot)}),$ (4)
where LN and $n_{u}$ denote layer normalization [42] for training stability and the number of sequential projection functions incorporating MLP for transformation and GeLU[43] for non-linearity, respectively. While our framework allows for repetition of the projection function to increase non-linearity, with 1 or 2 repetitions being sufficient (we use $n_{u}=2$, which aligns features as demonstrated in Fig. 1-(b)). Note that $\mathcal{U}^{s}$ is trained separately for each sensor based on reference query, which results in alignment of features from all sensors with respect to the fused feature as shown in Fig. 1-(b).
Cross-attention Across Sensors Along Patches with Post-feature Normalization (CASAP-PN). The patches $\mathbf{F}_{u}^{s}$ projected into the unified canonical space by UCP serve as keys ($K$) and values ($V$) for a trainable reference query $\mathbf{Q}_{ref}\!\!\in\!\mathbb{R}^{N_{q}\times C_{u}}$ (where $N_{q}$ is the number of queries $Q$), and we perform cross-attention across sensors along patches as:
$\mathbf{Q}_{ref,i}^{\prime}\!=\!\mathrm{CrossAttn}($ $Q\!=\!\mathbf{Q}_{ref},K\&V\!\in\!$ (5)
$\{\mathbf{F}_{u,i}^{S_{C}},\mathbf{F}_{u,i}^{S_{L}},\mathbf{F}_{u,i}^{S_{R}}\}),i\!=\!1:N_{p},$
where $\mathbf{Q}_{ref,i}^{\prime}$ is the output of the cross-attention applied across sensors for the $i$-th patch. Since $\mathbf{Q}_{ref}$ is trained primarily on features that are mostly available in the training data, it naturally develops high correlation (i.e., high attention scores) with patches from available sensors after the training. Consequently, during inference, $\mathbf{Q}_{ref,i}^{\prime}$ is predominantly composed of available $\mathbf{F}_{u}^{s}$. The number of heads in cross-attention is a hyper-parameter whose impact is analyzed in subsection 4.4.
Compared to ASF, cross-attention in SCF is performed across all positional-embedded patches with respect to object queries $\mathbf{Q}_{obj}$ in the transformer decoder head. This can be mathematically expressed as:
$\mathbf{Q}_{obj}^{\prime}\!=\!\mathrm{CrossAttn}^{(n_{td})}(Q\!=\!\mathbf{Q}_{obj},K\&V\!\in\!$ (6)
$\{\mathbf{F}_{pe,1}^{S_{C}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{C}}\!,\!\mathbf{F}_{pe,1}^{S_{L}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{L}}\!,\!\mathbf{F}_{pe,1}^{S_{R}}\!,\!...,\!\mathbf{F}_{pe,N_{p}}^{S_{R}}\!\}\!),$
where $\mathbf{F}_{pe}^{s}$ represents patches with positional-embedding, and $n_{td}$ (usually larger than 6 [24, 5]) is the number of stacked transformer decoders. Eq. 6 shows the cross-attention across all sensors and all patches (i.e., the K&V set contains $N_{s}N_{p}$ patches, resulting in $O(N_{s}^{2}N_{p}^{2})$ computational complexity). In contrast, in Eq. 5, the cross-attention is applied across sensors and along patches, which requires only $O(N_{s}^{2}N_{p})$ computational operations. This is a significant computational costs reduction as $N_{s}\ll N_{p}$. Moreover, as demonstrated in Fig. 3 and Tab. 1, ASF achieves better performance with only a single cross-attention layer than SCF utilizing stacked cross-attention layers (i.e., $n_{td}\!\geq\!6$).
Sequentially, ASF applies post-feature normalization (PN) $\mathcal{N}$ that has a similar structure to $\mathcal{U}^{s}(\cdot)$ with LN, to ensure that features can be processed consistently by the detection head regardless of the sensor combination. Therefore, PN enables camera, LiDAR, and 4D Radar features for the same object to be consistent. The set of patches $\mathcal{P}_{n}$ with PN is formulated as:
$\mathcal{P}_{n}\!=\!\{\mathbf{F}_{n,i}|\mathbf{F}_{n,i}\!=\!\mathcal{N}(\mathrm{LN}(\mathbf{Q}_{ref,i}^{\prime}))\!\in\!\mathbb{R}^{C_{u}}\!,i\!=\!1\!:\!N_{p}\}$ (7)
$\mathcal{N}(\cdot)\!=\!\mathrm{LN}(\mathrm{Proj^{(n_{n})}(\cdot)}),\mathrm{Proj(\cdot)\!=\!\mathrm{GeLU}(\mathrm{MLP}(\cdot)}),$ (8)
where $n_{n}$ denotes the number of sequential projections, with 1 or 2 being sufficient to increase non-linearity. Unlike Fig. 1-(b) where features from different sensors occupy distinct regions, Fig. 1-(c) illustrates how PN causes sensor features to converge into unified clusters.
Finally, to transform $\mathcal{P}_{n}$ back to the original BEV size ($H\!\times\!W$), we apply a reshape operation $\mathcal{T}_{n}(\cdot)$. Since the size of $\mathcal{P}_{n}$ is $C_{u}\!\times\!N_{p}\!=\!C_{u}\!\times\!N_{H}\!\times\!N_{W}\!=\!C_{u}\!\times\!(H/P_{H})\!\times\!(W/P_{W})$, the fused FM $\mathbf{FM}_{\mathrm{fused}}$ can be obtained as:
$\mathbf{FM}_{\mathrm{fused}}\!=\!\mathcal{T}_{n}(\mathcal{P}_{n}\!\in\!\mathbb{R}^{C_{u}\times N_{p}})\!\in\!\mathbb{R}^{C_{q}\times H\times W},$ (9)
where $C_{q}$ is the quotient of $C_{u}/(P_{H}\!\times\!P_{W})$ as we design $C_{u}\!=\!P_{H}\!\times\!P_{W}\!\times\!C_{q}$. Since the resulting channel dimension $C_{q}$ may be insufficient for containing feature representation due to reduced channel dimensions after reshaping, in the implementation, we increase the number of patches by a factor of $n_{p}$ (i.e., $N_{p}\!=\!N_{H}\!\times\!N_{W}\!\rightarrow\!N_{p}\!=\!n_{p}\!\times\!N_{H}\!\times\!N_{W}$). Consequently, the channel dimension of $\mathbf{FM}_{\mathrm{fused}}$ increases from $C_{q}$ to $n_{p}\!\times\!C_{q}$, and the impact of this modification is evaluated in subsection 4.4.
Methods Sensors IoU Metric Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
RTNH [12] (NeurIPS 2022) R 0.3 $AP_{BEV}$ 41.1 41.0 44.6 45.4 32.9 50.6 81.5 56.3
$AP_{3D}$ 37.4 37.6 42.0 41.2 29.2 49.1 63.9 43.1
0.5 $AP_{BEV}$ 36.0 35.8 41.9 44.8 30.2 34.5 63.9 55.1
$AP_{3D}$ 14.1 19.7 20.5 15.9 13.0 13.5 21.0 6.36
RTNH [12] (NeurIPS 2022) L 0.3 $AP_{BEV}$ 76.5 76.5 88.2 86.3 77.3 55.3 81.1 59.5
$AP_{3D}$ 72.7 73.1 76.5 84.8 64.5 53.4 80.3 52.9
0.5 $AP_{BEV}$ 66.3 65.4 87.4 83.8 73.7 48.8 78.5 48.1
$AP_{3D}$ 37.8 39.8 46.3 59.8 28.2 31.4 50.7 24.6
3D-LRF [2] (CVPR 2024) L+R 0.3 $AP_{BEV}$ 84.0 83.7 89.2 95.4 78.3 60.7 88.9 74.9
$AP_{3D}$ 74.8 81.2 87.2 86.1 73.8 49.5 87.9 67.2
0.5 $AP_{BEV}$ 73.6 72.3 88.4 86.6 76.6 47.5 79.6 64.1
$AP_{3D}$ 45.2 45.3 55.8 51.8 38.3 23.4 60.2 36.9
L4DR [34] (AAAI 2025) L+R 0.3 $AP_{BEV}$ 79.5 86.0 89.6 89.9 81.1 62.3 89.1 61.3
$AP_{3D}$ 78.0 77.7 80.0 88.6 79.2 60.1 78.9 51.9
0.5 $AP_{BEV}$ 77.5 76.8 88.6 89.7 78.2 59.3 80.9 53.8
$AP_{3D}$ 53.5 53.0 64.1 73.2 53.8 46.2 52.4 37.0
ASF (Proposed) L+R 0.3 $AP_{BEV}$ 88.6 88.1 90.3 99.0 89.1 80.4 89.4 78.7
$AP_{3D}$ 87.3 86.6 89.8 90.7 88.6 80.0 88.8 77.5
0.5 $AP_{BEV}$ 87.0 86.2 90.2 90.8 88.8 78.2 88.6 71.0
$AP_{3D}$ 72.9 64.6 86.6 79.6 73.4 70.0 77.6 66.7
C+L+R 0.3 $AP_{BEV}$ 88.6 88.2 90.2 98.9 89.0 80.4 89.2 78.4
$AP_{3D}$ 87.4 87.0 90.1 90.7 88.2 80.0 88.6 77.4
0.5 $AP_{BEV}$ 87.2 86.7 90.1 90.8 88.7 78.3 88.3 70.9
$AP_{3D}$ 73.6 71.8 87.0 79.4 73.0 67.5 78.0 66.4
Table 1: Performance comparison of 3D object detection on K-Radar [12] benchmark v1.0. C, L, and R represent Camera, LiDAR, and 4D Radar, respectively. The bold and underlined values indicate the best and the second-best, respectively. Note that the two ASF models for sensors L+R and C+L+R share the same neural network weights trained with C+L+R, which means that ASF for L+R represents a scenario where camera becomes unavailable.
3.3 Sensor Combination Loss (SCL)
Leveraging the consistent size of $\mathbf{FM}_{\mathrm{fused}}$ (which serves as the input to the detection head) regardless of sensor combinations, we propose an SCL that enables simultaneous optimization across multiple sensor configurations. The proposed SCL is formalized as:
$\mathcal{L}_{SCL}\!=\!\sum_{s\in\mathcal{SC}}(\mathcal{L}_{cls,s}\!+\!\mathcal{L}_{reg,s}),$ (10)
where $\mathcal{SC}$ represents the set of $7$ possible sensor combinations ($S_{C}$-only, $S_{L}$-only, $S_{R}$-only, $S_{L}\!+\!S_{R}$, $S_{C}\!+\!S_{R}$, $S_{C}\!+\!S_{L}$, and $S_{C}\!+\!S_{L}\!+\!S_{R}$), where $S_{C}$, $S_{L}$, and $S_{R}$ denote camera, LiDAR, and 4D Radar, respectively. $\mathcal{L}_{cls,s}$ and $\mathcal{L}_{reg,s}$ are the classification and regression losses for each sensor combination. SCL explicitly prepares for the potential sensor unavailability by optimizing across all sensor combinations in the training, enabling the model to recognize that certain sensors perform better than others (e.g., 4D Radar and LiDAR outperform cameras in adverse weather). As demonstrated in Tab. 4, SCL enhances the performance of the proposed ASF method when compared to ASF without SCL.
4 Experiments
4.1 Experimental Setup
Dataset and Metrics. K-Radar [12] is a large-scale AD dataset with a broad range of conditions including time (day, night), weather (normal, rain, fog, snow, sleet), road types (urban, highway, mountain), and sensors (4D Radar, LiDAR, camera, GPS). Notably, K-Radar is the only dataset with data captured in adverse weather conditions.
To compare ASF with the existing SOTA methods, we use two official variants of K-Radar settings. Following the protocol in [12] and [2, 34], we evaluate ASF and SOTA methods using benchmark v1.0, which focuses on the ‘Sedan’ class with label v1.0 and covers a region close to the driving corridor: [0m, 72m] for X-axis, [-6.4m, 6.4m] for Y-axis, and [-2m, 6m] for Z-axis. Additionally, we present ablation studies and qualitative results on benchmark v2.0, which analyzes a wider area ([0m, 72m] for X-axis, [-16m, 16m] for Y-axis, and [-2m, 7.6m] for Z-axis) and includes both ‘Sedan’ and ‘Bus or Truck’ classes (two dominant classes in [12]). For evaluation metrics, we use: $AP_{3D}$ and $AP_{BEV}$ at IoU thresholds of $0.3$ (following the protocol in [12]) and $0.5$ (for a strict assessment). Additionally, we report the required VRAM and FPS (frames per second) based on the same hardware setup.
Implementation Details. We implement the ASF on a single RTX3090 GPU with 24GB VRAM. ASF is trained for 11 epochs using AdamW [44] optimizer with a learning rate $0.001$ and a batch size $2$. The voxel size for the fused FM is set to $0.4$m, consistent with [12].
Figure 3: Qualitative results of ASF for various sensor combinations. We show results for normal and adverse weather conditions in (a-i) and (j-r), respectively, where employed sensors are noted in the top-left corner (C: Camera, L: LiDAR, R: 4D Radar, C*: damaged camera). Each subplot visualizes front-view camera image, LiDAR point cloud, 4D Radar tensor, and a sensor attention map (SAM) showing attention score distribution from cross-attention in CASAP-PN. In the SAMs, red, green, and blue represent attention scores for Camera, LiDAR, and 4D Radar, respectively. For example, a predominantly blue SAM indicates that 4D Radar received the highest attention scores in the scene. The bottom-left corner of each subplot shows the proportion of attention scores in colored percentages (C/L/R[%]). Note that predictions are visualized on all sensor data, even when a sensor is not employed for detection (e.g., predictions from L+R are also visualized on the camera image).
4.2 Comparison of ASF to SOTA methods
Following the benchmark v1.0 of the K-Radar [12], we compare the proposed ASF with SOTA methods including 3D-LRF [2] and L4DR [34], and we use RTNH [12] for single-sensor performance. In addition to the detection performance comparison with DCF methods [2, 34], we evaluate computational efficiency against DPFT [3] which is the only open-sourced SCF method available for K-Radar.
Methods Sensors VRAM [GB] FPS [Hz]
3D-LRF [2] L+R 1.2 5.04
DPFT [3] C+R 4.0 10.8
ASF L+R 1.5 20.5
(Proposed) C+L+R 1.6 13.5
Table 2: Comparison of VRAM and FPS evaluated on the K-Radar benchmark v1.0. The bold and underlined values indicate the best and the second-best, respectively.
Class Sensors Total Normal Ove. Sleet H.s.
Sedan R 47.3 40.7 58.8 45.9 56.5
L 73.0 73.0 86.1 64.9 54.5
C 14.8 14.9 7.71 - -
C 3.71 3.72 3.17 - -
L+R 77.3 77.7 87.3 74.4 65.4
C+R 52.7 49.1 62.4 46.0 57.2
C+L 76.4 78.3 86.5 64.2 57.1
C+L+R 79.3 78.8 87.6 74.2 65.8
C+L+R 77.6 78.2 87.7 74.4 65.4
C+L+R 58.9 58.8 66.6 52.3 58.2
Bus or Truck R 34.2 22.9 40.9 21.1 51.2
L 54.9 53.7 74.8 69.1 37.8
C 9.59 9.02 17.2 - -
C 3.65 3.66 0.00 - -
L+R 59.9 52.5 71.6 68.2 68.9
C+R 36.2 24.4 41.4 23.4 56.5
C+L 53.0 49.1 60.1 72.1 39.6
C+L+R 60.4 52.7 77.4 69.2 68.9
C+L+R 60.1 52.1 72.0 70.9 69.1
C+L+R 40.0 31.5 38.2 28.9 54.8
Table 3: Performance comparison of ASF under various sensor combinations on K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.3 for ‘Sedan’ and ‘Bus or Truck’ classes. ‘Ove.’ and ‘H.s.’ refer to ‘Overcast’ and ‘Heavy Snow’, respectively. * denotes the sensor unavailability, as shown in Fig. 3. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained for R+L+C. Performance under additional weather conditions (Fog, Rain, and Light Snow) and with other evaluation metric $AP_{BEV}$ is provided in Appendix C.
Detection Performance. As shown in Tab. 1, ASF significantly outperforms SOTA methods for various weather conditions. Compared to the previous SOTA method L4DR [34], ASF shows substantial improvements of 9.7% in $AP_{BEV}$ (i.e., ASF 87.2% vs. L4DR 77.5%) and 20.1% in $AP_{3D}$ (i.e., ASF 73.6% vs. L4DR 53.5%) at IoU=0.5. These improvements are particularly significant for IoU=0.5 (a strict evaluation metric), highlighting ASF’s superior detection performance. The performance advantages are consistent across all weather conditions, with especially notable improvements in challenging sensor degradation scenarios such as sleet (67.5% vs. 46.2% $AP_{3D}$ at IoU=0.5) and heavy snow (66.4% vs. 37.0% $AP_{3D}$ at IoU=0.5).
In addition, we note both ASF (L+R and C+L+R) use the same neural network weights, yet maintain comparable high performance levels. This means that even when camera images become severely degraded, the system gracefully maintains detection performance with the remaining sensors (L+R). This capability is crucial for reliable AD that should be able to cope with unexpected sensor failure. As in Tab. 1, ASF using only R+L achieves 87.0% $AP_{BEV}$ and 72.9% $AP_{3D}$ at IoU=0.5, which is very close to the performance of the full sensor suite (87.2% $AP_{BEV}$ and 73.6% $AP_{3D}$).
Computational Efficiency. The computational efficiency of ASF is another important requirement for real-time AD. As shown in Tab. 2, ASF achieves significant fast processing speed while maintaining low memory requirements. ASF with LiDAR and 4D Radar operates at 20.5 Hz, which is about 4 times (i.e., 4$\times$) faster than 3D-LRF (5.04 Hz) when using the same sensors. Even when all three sensors (C+L+R) are used, ASF maintains 13.5 Hz that exceeds the typical requirement of 10 Hz for AD [8]. This efficiency stems from the proposed CASAP-PN, where cross-attention is applied across sensors along patches rather than across all sensors and all patches as in SCF methods. Additionally, our unified canonical projection eliminates the requirement for computationally expensive positional embedding. The memory footprint of ASF remains remarkably low (1.5-1.6 GB), which is comparable to 3D-LRF (1.2 GB) and much lower than DPFT (4.0 GB).
Exp. $P$ $C_{u}$ $n_{p}$ $n_{h}$ SCL $AP_{3D}^{Sedan}$ $AP_{3D}^{Bus}$
(a) 5 512 1 8 76.1 45.7
(b) 5 512 4 8 76.4 47.4
(c) 2 512 4 8 77.2 49.7
(d) 2 256 8 8 77.6 57.9
(e) 2 256 8 8 $\checkmark$ 79.3 58.2
(f) 2 256 8 16 77.5 60.2
(g) 2 256 8 16 $\checkmark$ 79.3 60.4
Table 4: Ablation study of ASF. ASF performance for different components and parameters: $P$ (patch size), $C_{u}$ (channel dimension in unified canonical space), $n_{p}$ (number of patches multiplier), $n_{h}$ (number of heads in CASAP-PN) and SCL, using $AP_{3D}$ at IoU=0.3 for both ‘Sedan’ and ‘Bus or Truck’ on the K-Radar benchmark v2.0.
4.3 Addressing Sensor Degradation and Failure
A key advantage of ASF is the robust performance under sensor degradation or failure. As shown in Tab. 3 and Fig. 3, ASF dynamically adapts to different sensor combinations without retraining. Under normal conditions (Fig. 3-(a-i)), ASF effectively utilizes all available sensors with attention weights distributed according to each sensor’s reliability. However, ASF’s true value emerges in challenging scenarios. In adverse weather (Fig. 3-(j-r)), camera and LiDAR measurements are significantly degraded or disappear completely. In these critical situations, ASF automatically redistributes attention toward the more reliable 4D Radar, as evidenced by the predominant blue coloration in the sensor attention maps (SAMs) and corresponding attention percentages. Even with damaged sensors (denoted by * in Tab. 3 and Fig. 3), ASF maintains near-optimal performance. For example, with a damaged camera (C*), C*+L+R shows 77.6% AP3D, which is only 1.7% lower than with fully functional sensors (79.3%). This robustness stems from the unified canonical projection (which creates a common feature space) and the cross-attention mechanism (which implicitly estimates sensor reliability).
The qualitative results in Fig. 3 demonstrate that in adverse weather, when LiDAR measurements disappear and camera visibility severely degrades, reliable object detection is only possible with active 4D Radar and ASF is fully using 4D Radar. Note that all results shown are from the same ASF model with identical weights, illustrating how ASF dynamically adjusts attention to maintain detection performance across varying sensor availabilities.
4.4 Ablation Studies
Tab. 4 presents ablation studies of key components of ASF. We analyze five factors: patch size ($P$), channel dimension ($C_{u}$), patches multiplier ($n_{p}$), number of attention heads ($n_{h}$), and sensor combination loss (SCL). Comparing experiments (a)-(c), we observe that smaller patch sizes ($P=2$) improve performance by allowing more fine-grained feature extraction. Experiment (d) shows that reducing channel dimension from $C_{u}=512$ to $C_{u}=256$ while increasing the patches multiplier ($n_{p}$) from 4 to 8 maintains or improves the performance, suggesting a better balance between channel depth and spatial representation. Comparing experiments (d) and (f), we find that the number of attention heads impacts classes differently - more heads ($n_{h}=16$) significantly benefit the ’Bus or Truck’ class while showing minimal effect on the ’Sedan’ class. Experiments (e) and (g) demonstrate that incorporating SCL consistently improves performance across configurations, confirming that explicitly training for multiple sensor combinations enhances robustness during inference. The optimal configuration (g) uses $P\!=\!2$, $C_{u}\!=\!256$, $n_{p}\!=\!8$, $n_{h}\!=\!16$, and includes SCL.
5 Conclusion
This paper introduces availability-aware sensor fusion (ASF), which addresses sensor availability challenges in AD by transforming features into a unified canonical space through UCP and CASAP-PN. Our approach maintains computational efficiency ($O(N_{s}^{2}N_{p})$) while providing robust fusion for sensor degradation or failure. The proposed sensor combination loss further enhances robustness by optimizing across all possible sensor combinations. Experiments on the K-Radar dataset demonstrate significant improvements over SOTA methods (9.7% in $AP_{BEV}$ and 20.1% in $AP_{3D}$ at IoU=0.5), with consistent performance across various weather conditions and sensor combinations.
Limitations. Despite ASF’s strong performance, the camera network’s capabilities remain a limitation. As shown in Fig. 3-(g) and (p), camera-based object detection is less precise, particularly in adverse weather. In Fig. 1-(c), while LiDAR and 4D Radar features are well integrated, camera features remain more separated in feature space. Enhancing the camera backbone could further boost system performance, especially in favorable weather conditions, where visual information is valuable.
Appendix. The appendix is organized as follows. Section A provides a detailed explanation of the sensor fusion framework, including specific implementations of each component. Section B presents additional t-SNE visualizations that further demonstrate the effectiveness of the proposed method. Section C includes additional performance results across various weather conditions and sensor combinations.
Appendix A Details of Sensor Fusion Framework
Figure 4: Overall sensor fusion framework for camera, LiDAR, and 4D Radar. ‘FM’ denotes ‘feature map’.
Figure 5: Visualization of feature representations through t-SNE [1] at different stages of the proposed ASF for the ‘Bus or Truck’ class. ‘UCP’ and ‘CASAP-PN’ denote unified canonical projection and cross-attention across sensors along patches with post-feature normalization, which are the two main components of the proposed ASF.
As shown in Fig. 4, the sensor fusion framework is divided into three stages: (1) sensor-specific encoders [45, 35, 39, 46, 12, 15] that extract feature maps (FMs) from different types of multi-modal sensors (e.g., camera, LiDAR, and 4D Radar), (2) a fusion method that integrates these FMs, and (3) a task-dependent detection head. Stage (1) has evolved independently for each sensor [47, 48, 12], while stage (3) employs a task-dependent method (e.g., YOLO head [49] for object detection) regardless of sensor type. Therefore, in this work, we propose a fusion method focusing on stage (2) rather than stages (1) and (3). Details of both existing and proposed fusion methods are illustrated in Fig. 2. In this section, we describe both the sensor-specific encoders and the detection head of the proposed availability-aware sensor fusion (ASF).
A.1 Sensor-specific Encoders
Sensor-specific encoders extract feature maps (FMs) from each sensor’s data. Because different sensors provide data in different modalities—cameras produce 2D RGB images, LiDAR generates 3D point clouds, and 4D Radar outputs low-resolution tensors with power values—research on these encoders has evolved independently for each sensor type.
For ASF, we adopt established backbone architectures for each sensor and pre-train these backbones before ASF training. Each encoder is trained for 11 epochs following the K-Radar [12] training protocol. During ASF training, we freeze all sensor-specific encoders to ensure that only the ASF components (UCP and CASAP-PN) are learned. Details for each encoder are as follows:
Camera. Camera-based sensor encoders for autonomous driving typically require transforming front-view (FV) images into a bird’s-eye-view (BEV) format for 3D object detection. Current approaches can be broadly categorized into two methods: (1) learnable BEV transformations such as Lift-Splat-Shoot (LSS) [35, 18, 17], or (2) maintaining FV while using transformer decoders [24] that incorporate depth information as positional embeddings in BEV space. Although methods like PETR [24] have improved 3D object detection by injecting depth embeddings, they inevitably increase computational complexity due to the additional depth dimension. These two approaches form the foundation of camera-based fusion [20, 19, 18, 17, 5], commonly seen in deeply coupled fusion (DCF) and sensor-wise cross-attention fusion (SCF).
To develop a computationally efficient method, we opt for the LSS approach and adopt BEVDepth [38], which supervises the LSS depth estimation using LiDAR point clouds. For encoding images, we use the Swin-S model of SwinTransformer as the 2D backbone, achieving 83% top-1 accuracy on ImageNet [50]. We resize K-Radar’s front camera images to $704\!\times\!256$ pixels for encoding. The encoded 2D feature maps are projected into BEV feature maps via LSS with a grid size of 0.2 m. To match the 0.4 m grid size used in K-Radar, we add a convolutional layer with stride 2. During training, we employ the depth information projected from LiDAR point clouds onto the image and, inspired by [51], we use the LiDAR backbone’s (SECOND) BEV feature maps for distillation. As shown in Fig. 3, camera images have limited utility in adverse weather, so we train the camera encoder only on normal or overcast scenes (sequences 1–20) from K-Radar.
A limitation of this approach is that promising stereo or multi-view detection models cannot be fully leveraged, because the K-Radar dataset [12]—the only one offering camera, LiDAR, and 4D Radar data in adverse weather—provides object annotations only for the front view. Moreover, unlike the KITTI dataset with its longer stereo baseline of 0.54 m, K-Radar uses waterproof ZED cameras with a baseline of just 0.12 m, limiting stereo-based detection for distant objects. Consequently, we employ a monocular camera method despite inherent performance constraints.
LiDAR. LiDAR provides precise depth information in the form of dense point clouds, often used as a reference for labeling other sensors [12, 52]. We adopt the SECOND [39] model as the LiDAR encoder, which is open-source in K-Radar and widely used in various benchmarks. Leveraging LiDAR’s high spatial resolution, we set the voxel size to 0.05 m and apply multiple layers of sparse convolution to generate BEV feature maps with a 0.4 m grid size.
4D Radar. Unlike LiDAR, which uses Time-of-Flight (ToF) measurement, Radar detects spatial information through signal processing, which can introduce sidelobe noise. Traditional 3D Radar typically filters out noisy measurements to produce a sparse point cloud focusing on object presence [11]. In contrast, 4D Radar captures additional elevation information, yielding higher-density point clouds [53]. This property is exemplified by K-Radar [12, 14]. We adopt RTNH [12], the baseline provided in K-Radar, as our 4D Radar encoder, following the original network settings and design.
A.2 Detection Head
Following the K-Radar baseline, we use an SSD detection head [40, 41] for the 3D object detection task. Our implementation uses two anchor boxes per object class, oriented at $0^{\circ}$ and $90^{\circ}$. Each anchor box is parametrized by eight values: the 3D center coordinates ($x_{c}$, $y_{c}$, $z_{c}$), 3D size ($x_{l}$, $y_{l}$, $z_{l}$), and orientation ($\cos(\text{yaw})$, $\sin(\text{yaw})$), as in [41].
To address the severe class imbalance in object detection (negative samples far outnumber positives), we use the focal loss [54] for classification. For bounding box regression, we use smooth L1 loss to measure differences between predictions and ground truth, improving training stability. During inference, we select the class with the highest logit score for each proposal, then apply non-maximum suppression to remove redundant overlapping detections, producing the final predictions.
Figure 6: Qualitative results of ASF with C+L+R (C: Camera, L*: damaged LiDAR, R: 4D Radar). We show results under (a–b) normal and (c–d) adverse weather conditions. Each grid shows the camera image, LiDAR point cloud, 4D Radar tensor, and a sensor attention map (SAM) illustrating attention scores from CASAP-PN (except for (a) and (c), which only show sensor data for ground truth visualization). In the SAM, red, green, and blue represent attention to Camera (C), LiDAR (L), and 4D Radar (R), respectively. For example, a predominantly blue SAM indicates that 4D Radar has the highest attention in that scene. The bottom-left corner of each grid shows attention score proportions (C/L/R[%]). Yellow and red boxes denote predictions and ground truth, respectively.
Appendix B t-SNE Visualizations
B.1 Additional t-SNE Visualizations
This section provides additional feature visualizations to complement those in Fig. 1, which only shows results for the ‘Sedan’ class. Fig. 5 presents t-SNE [1] visualizations of the ‘Bus or Truck’ class at different stages of the ASF. These visualizations further demonstrate how ASF unifies feature representations from different sensors.
Consistent with Fig. 1, we observe three main trends in the t-SNE visualizations. First, the initial features from the sensor-specific encoders are scattered without clear alignment across different sensors. Next, after the unified canonical projection (UCP), sensor-specific features become better aligned with the fused features. Finally, following CASAP-PN, features from available sensors form cohesive clusters (in this case, LiDAR and 4D Radar).
These findings underscore the effectiveness of ASF in creating a unified canonical space for multi-sensor features. Note that in K-Radar, the ‘Sedan’ class has 37,213 instances, whereas the ‘Bus or Truck’ class has 7,449 instances, leading to somewhat sparser data in Fig. 5 compared to Fig. 1.
B.2 Details of t-SNE Visualizations
To visualize object-specific features in each sensor’s feature map (FM), we use RoIAlign [55] to crop features within bounding boxes. RoIAlign extracts fixed-size feature maps, making it suitable for t-SNE (which requires consistent dimensions). We then visualize these cropped features in the unified canonical space. For those extracted before projection (where channel dimensions may differ), we downsample as needed to ensure consistent feature dimensions.
Appendix C Additional Detection Results
This section presents additional detection results for ASF. First, we provide qualitative examples where the LiDAR sensor is damaged. Then, we show quantitative performance, including $AP_{3D}$ and $AP_{BEV}$ for both ‘Sedan’ and ‘Bus or Truck’ at IoU=0.3 and 0.5 on K-Radar [12] benchmark v2.0. These complement Tab. 3, which only reports $AP_{3D}$ at IoU=0.3.
C.1 Additional Qualitative Results on Damaged LiDAR
When LiDAR surfaces are damaged or contaminated, LiDAR measurements can be lost [56]. In adverse conditions such as heavy snow, accumulation on the LiDAR sensor can block forward measurements, as illustrated by the red dashed line in Fig. 6-(d), even though vehicles are present. We also simulate total loss of forward LiDAR returns from surface damage (e.g., broken sensor), as seen in Fig. 6-(b). Despite missing LiDAR data, ASF still detects all objects using only camera and 4D Radar inputs. As reported in Tab. 3, ASF with C+L+R often performs similarly or even better than C+R alone. The sensor attention map (SAM) in Fig. 6-(b) confirms the increased reliance on camera or 4D Radar when LiDAR is unavailable.
Class Sensors Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
Sedan R 47.3 40.7 58.8 84.4 41.0 45.9 66.1 56.5
L 73.0 73.0 86.1 92.9 64.3 64.9 86.0 54.5
C 14.8 14.9 7.7 - - - - -
C 3.7 3.7 3.2 - - - - -
L+R 77.3 77.7 87.3 95.5 71.0 74.4 91.0 65.4
C+R 52.7 49.1 62.4 84.6 44.6 46.0 66.5 57.2
C+L 76.4 78.3 86.5 93.4 69.4 64.2 87.3 57.1
C+L+R 79.3 78.8 86.1 93.7 72.9 74.2 91.2 65.8
C+L+R 77.6 78.2 87.7 93.8 71.3 74.4 91.1 65.4
C+L+R 58.9 59.2 64.0 92.4 47.7 65.4 66.5 58.2
Bus or Truck R 34.2 22.9 41.0 - 0.2 21.1 83.5 51.2
L 54.9 53.7 74.8 - 5.1 69.1 85.2 37.8
C 9.6 9.0 17.2 - - - - -
C 3.7 3.7 0.0 - - - - -
L+R 59.9 52.5 71.6 - 6.8 68.2 88.1 68.9
C+R 36.2 24.4 41.4 - 0.3 23.4 82.1 56.5
C+L 53.0 49.1 60.1 - 4.4 72.1 85.1 39.6
C+L+R 60.4 52.7 77.4 - 8.0 69.2 87.9 68.9
C+L+R 60.1 52.1 72.0 - 8.4 70.9 87.9 69.1
C+L+R 40.0 31.5 38.2 - 0.4 28.9 81.0 54.8
Table 5: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.3 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
Class Sensors Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
Sedan R 55.9 48.7 70.1 92.0 50.8 52.6 76.8 61.1
L 76.7 78.1 93.8 93.6 70.0 67.8 88.4 58.2
C 17.6 17.7 8.9 - - - - -
C 5.3 5.3 3.5 - - - - -
L+R 81.8 80.7 94.3 98.1 75.7 81.0 93.6 69.6
C+R 58.7 54.7 71.1 91.9 52.7 52.4 74.6 61.3
C+L 79.3 81.0 94.9 95.1 72.5 68.3 89.7 59.5
C+L+R 82.5 81.8 94.9 98.0 76.2 81.0 93.8 69.5
C+L+R 82.1 81.1 94.7 98.1 76.0 80.2 93.8 69.4
C+L+R 64.5 64.6 76.1 92.5 57.2 56.9 75.9 60.6
Bus or Truck R 43.5 28.4 42.9 - 0.3 23.7 92.5 77.2
L 63.1 57.0 75.5 - 6.2 72.3 88.7 61.3
C 11.4 11.2 18.1 - - - - -
C 3.9 4.0 0.0 - - - - -
L+R 69.5 58.3 75.2 - 7.3 71.3 96.2 89.7
C+R 44.6 29.2 43.4 - 0.3 26.0 91.1 76.9
C+L 61.8 53.1 68.0 - 5.3 75.5 90.0 64.7
C+L+R 70.2 58.3 81.5 - 8.0 72.2 96.1 89.6
C+L+R 70.0 58.2 75.7 - 9.2 73.8 96.4 88.4
C+L+R 48.0 36.2 42.6 - 0.4 33.6 91.0 77.2
Table 6: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{BEV}$ at IoU = 0.3 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
Class Sensors Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
Sedan R 20.4 13.2 26.6 59.5 18.1 20.9 28.7 30.7
L 41.1 38.5 43.1 67.8 37.1 34.1 52.4 35.4
C 5.9 5.9 2.8 - - - - -
C 1.0 1.0 0.8 - - - - -
L+R 52.1 49.0 56.9 81.4 48.2 45.1 59.6 50.1
C+R 27.5 22.6 29.8 66.2 23.1 27.2 30.2 39.3
C+L 52.3 52.3 50.0 79.4 45.7 40.1 65.9 42.7
C+L+R 52.9 50.5 55.8 81.5 49.1 44.6 60.1 50.0
C+L+R 52.3 49.3 56.8 79.6 48.9 45.0 60.3 50.0
C+L+R 33.7 29.9 32.3 67.5 29.0 29.1 37.6 38.7
Bus or Truck R 16.5 8.5 22.6 - 0.1 8.7 65.3 24.6
L 29.1 22.1 38.0 - 1.1 44.2 66.2 20.5
C 2.6 2.5 6.3 - - - - -
C 2.6 2.6 0.0 - - - - -
L+R 31.7 24.2 52.6 - 2.8 44.4 70.9 33.2
C+R 18.1 9.7 18.3 - 0.2 8.7 65.1 28.0
C+L 31.2 22.5 33.6 - 0.9 51.4 69.6 21.2
C+L+R 31.6 23.7 47.0 - 2.5 45.0 72.3 33.0
C+L+R 32.1 24.5 48.2 - 2.3 44.8 70.6 33.3
C+L+R 19.2 13.5 15.6 - 0.0 9.8 62.5 27.9
Table 7: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{3D}$ at IoU = 0.5 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
Class Sensors Total Normal Overcast Fog Rain Sleet Light snow Heavy snow
Sedan R 40.8 30.0 54.7 85.6 36.9 39.6 65.2 48.4
L 66.8 65.7 79.0 90.7 60.1 54.6 83.4 52.7
C 8.8 8.9 4.3 - - - - -
C 2.1 2.1 1.1 - - - - -
L+R 74.0 72.6 84.4 96.1 67.3 67.4 90.8 63.4
C+R 45.8 39.7 57.2 87.4 40.7 41.6 66.7 53.9
C+L 72.3 72.8 87.2 93.3 62.5 55.4 87.2 55.9
C+L+R 74.5 74.8 85.1 96.2 67.0 67.5 90.8 63.5
C+L+R 74.1 72.9 85.0 96.1 66.9 66.7 90.7 63.2
C+L+R 53.1 54.5 57.0 61.5 50.2 50.3 68.5 55.2
Bus or Truck R 23.8 14.2 30.0 - 0.2 12.7 87.3 64.3
L 45.8 35.7 61.1 - 2.3 58.2 81.9 53.3
C 11.4 11.2 18.1 - - - - -
C 2.7 2.8 0.0 - - - - -
L+R 53.1 37.5 59.8 - 6.1 56.6 88.2 80.1
C+R 31.9 14.4 25.9 - 0.2 13.4 90.1 64.7
C+L 47.3 32.4 40.7 - 3.3 63.1 85.3 55.6
C+L+R 53.9 36.5 54.7 - 7.9 59.3 90.4 80.4
C+L+R 53.6 36.1 57.5 - 8.0 59.7 90.8 80.1
C+L+R 33.2 20.9 24.7 - 0.2 18.3 88.0 61.4
Table 8: Performance comparison of ASF under various sensor combinations evaluated on the K-Radar [12] benchmark v2.0. We indicate the employed sensors (C: Camera, L: LiDAR, R: 4D Radar) and report $AP_{BEV}$ at IoU = 0.5 under various weather conditions for the ‘Sedan’ and ‘Bus or Truck’ classes. The bold and underlined values indicate the best and the second-best, respectively. All ten ASF models share the same neural network weights trained with R+L+C.
C.2 Additional Quantitative Results
Tables 5, 6, 7, and 8 present comprehensive evaluations of ASF performance across various sensor configurations and weather conditions. Results demonstrate that multi-sensor fusion approaches consistently outperform single-sensor methods. The C+L+R configuration achieves the highest performance in most scenarios for both ‘Sedan’ and ‘Bus or Truck’ classes, with C*+L+R often ranking second, which indicates that a damaged camera does not significantly affect performance.
Performance analysis across weather conditions reveals that radar-inclusive combinations (C+L+R, L+R) exhibit superior performance in adverse weather (demonstrating 4D Radar’s effectiveness under challenging environmental conditions), particularly in fog and snow. While camera-LiDAR fusion (C+L) occasionally outperforms in normal and overcast conditions, its performance deteriorates significantly in precipitation.